# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ 5 Charged particles in a magnetic field
The statement that a fractionally quantized Hall conductivity implies the emergence of Abelian anyon excitations with fractional charge is well-known. When restated in terms of the higher-form symmetry of the low-energy theory, the conclusion of Sec. 4 might look strange. Where did the one-form symmetry of the low-energy theory come from? An obvious possible answer is that it can be an emergent symmetry, i.e., a symmetry of the IR theory without a precursor in the UV. Although this is certainly a possibility, in this section, we will demonstrate that in the case of particles in a uniform magnetic field, the IR one-form symmetry and its anomaly started their lives in the UV theory as ordinary space translations.
@@ 5.1 The UV theory
[…]
@@
The parameters $\sigma_{i}$ correspond to $\delta A_{i}={\sigma_{i}/L_{i}}$. Changing the origin of $x$ or $y$ shifts $\sigma_{y}$ and $\sigma_{x}$ Hence, these constants signal the fact that the classical $\mathrm{U}(1)$ translation symmetry around every direction is only $\mathbb{Z}_{\cal K}$, rather than $\mathrm{U}(1)$. They are generated by
@@
In the first equation in (5.8), $\sigma_{i}$ are related to a background $\delta A_{i}^{(0)}$ and can be thought of as due to topological defects associated with the global $\mathrm{U}(1)$ symmetry. (Standard flux threading discussions analogous to those in Sec. 2.2 and its counterpart in Sec. 4 shift these values.) The analysis of such defects leads to a mixed anomaly between the discrete ${\mathbb{Z}}_{\cal K}$ translations and $\mathrm{U}(1)$, which in turn leads to the famous filling constraints [83, 84, 85, 86]. Such a change of the background field by $\sigma_{i}$ with vanishing field strength is often referred to as “flux insertion,” and it is done when the total field strength vanishes. Here, the background magnetic field is non-zero, but we can still deform the theory with $\delta A$ with vanishing field strength and use the standard techniques of such defects.
The second equation in (5.8) means that this discrete translation symmetry is realized projectively. This can be interpreted as anomalies in the discrete translation symmetry and the global $\mathrm{U}(1)$ symmetry. [^23] See App. A in [87], for a pedagogical discussion of the well-known Landau levels from the perspective of ’t Hooft and Adler-Bell-Jackiw anomalies. This interpretation is significant because it leads to ’t Hooft anomaly matching conditions. In particular, the same anomalous phases should exist in the IR theory. We will use this constraint to classify the possible gapped phases.
@@ 5.2 From the UV to the IR
Let us start with the third question. Most of the symmetry (5.8) is anomaly-free, and therefore it can act trivially in the IR. Let the generators $T^{i}$ flow in the IR to $T_{\rm IR}^{i}$. Then, the elements $(T_{\rm IR}^{i})^{q}$ do not participate in the anomaly and they can act trivially; i.e., act as central elements in the IR theory. We will allow a more general situation depending on another integer of order one $\ell$ and define [^24] This matches with the notation used in Sections 4 and 7. Comparing with the discussions in App. B, here we have $u=1$.
@@
${q\over 4\pi}\hat{b}\wedge d\hat{b}-{1\over 2\pi}\delta A\wedge d\hat{b}\,$ (5.14)
@@
where isubscriptcontour-$\oint_{i}$ is a contour integral around the $i$ direction. (For odd $q$ we also need a transparent fermion line on the right-hand side.) This is consistent with the fact that this theory has $p=r=1$.
@@
Finally, given that the IR version of the translation generators $T_{\rm IR}^{i}$ act in the IR as one-form symmetry operators, we can tie this system to the flux threading argument in Sections 2.2 and 4. We identify $T_{\rm IR}^{i}$ with the one-form symmetry generated by the vison $v$. Then, the first equation in (5.11) is the statement that $v$ carries charge $r/s=p/q$, the second equation in (5.11) is the statement that $B(v,v)=r/s=p/q$, and equation (5.10), which states which subgroup of the translation group is realized trivially in the IR, determines the power of the vison that is trivial. In other words, we have learned that the UV precursor of the vison $v$ is a UV translation symmetry generator.
@@ 6 Connection between translation and flux threading
We expect a gapped ground state to be described by a TQFT, which might be trivial. Then, from the Lieb-Schulz-Mattis-Oshikawa-Hastings theorem [90, 83, 91, 92], we know that for fractional $\nu_{l}$ gapped ground states must have nontrivial topological order. In particular, they must have an Abelian background anyon, $a_{b}$, with charge $Q(a_{b})=\nu_{l}$. This is because translation acts projectively on every anyon $a$ as $(\hat{T}^{y})^{-1}(\hat{T}^{x})^{-1}\hat{T}^{y}\hat{T}^{x}[a]=e^{2\pi i\varphi(a)}[a]$. [^27] Here $[a]$ refers to the state of a single anyon. In the continuum theory, this is a defect line of $a$ stretching along the time direction. In other words, the action of translation is given by a phase, $\varphi$, that respects fusion. (We restrict attention to phases where the action of translation does not permute anyons.) The results of [57], summarized in App. C, thus tell us that we can write $(\hat{T}^{y})^{-1}(\hat{T}^{x})^{-1}\hat{T}^{y}\hat{T}^{x}[a]=e^{2\pi iB(a,a_{b})}[a]$ for some Abelian $a_{b}$. The projective action of translation thus means we can think of the background as having $a_{b}$ in each unit cell [93, 94, 95, 84, 96]. As such $a_{b}$ is often called the “background anyon.”
@@
Now, consider the action of a unit translation on the ground state. For concreteness, consider a system with $N_{x}\times N_{y}$ unit cells placed on a torus. We label the unit cells by $\vec{R}=m_{x}\hat{x}+m_{y}\hat{y}$, $m_{i}=1,.....,N_{i}$. Acting with, say, $\hat{T}^{x}$, shifts $a_{b}(\vec{R})$ to $a_{b}(\vec{R}+\hat{x})$. Equivalently, we can think of this process as creating a separate $a_{b}-\bar{a_{b}}$ pair along each of the $N_{y}$ chains labeled by distinct $m_{y}$, moving each $a_{b}$ around its chain and annihilating with $\bar{a_{b}}$. In other words, $\hat{T}^{x}$ maps, in the IR theory, to a line operator of $a_{b}$ on each of the $N_{y}$ chains. Now consider acting with $\hat{T}^{x}$ in a ground state, where there is a vison line stretched along the $y$-cycle. Such a vison line can be created by “threading a $2\pi$ flux of the background $\mathrm{U}(1)$ gauge field” (i.e., creating a holonomy of the background $\mathrm{U}(1)$ gauge field). More precisely, starting with any ground state, we first apply a vison line operator along the $y$-cycle, and then act on the state with $\hat{T}^{x}$, and compare the result with the one obtained from acting in the opposite order. Due to the mutual braiding between $a_{b}$ and $v$, the $\hat{T}^{x}$ action now produces a phase $2\pi\nu_{l}N_{y}$. We show a cartoon of this in Fig. LABEL:fig:translation. This is exactly the answer one would infer from the Oshikawa adiabatic flux-threading argument applied to this system [96].
@@
Then, from our discussion earlier, the action of $\hat{T}^{x}$ corresponds to a line operator of $v$ around the $x$-cycle. Since the unit cell has length $L_{y}/n_{x}$ along the $y$-direction, there are $n_{x}$ defect loops parallel to each other. We see that the action of $\hat{T}^{x}$ corresponds, in the IR theory, to a line operator of $v^{n_{x}}$ along the $x$-cycle. This identification does not depend on the strength of the perturbing potential, and hence holds even in the limit that the potential is turned off. Repeating the same argument for the $y$-direction, we see that $\hat{T}^{y}$ corresponds in the IR to a line defect of $v^{n_{y}}$ along the $y$-cycle.
In Sec. 5, we discussed the magnetic translation symmetry $\mathbb{Z}_{\cal K}\times\mathbb{Z}_{\cal K}$ of particles in a homogeneous magnetic field with its anomaly ${\cal N}\over{\cal K}$. The relation to the $\mathbb{Z}_{n_{x}}\times\mathbb{Z}_{n_{y}}$ translation symmetry discussed here is that the latter is an anomaly-free subgroup of the former. For example, we can specialize to $n_{x}=1,n_{y}={\cal K}$, which leads to the conclusion that $\hat{T}^{x}=T^{x}$ creates a line defect of $v$ along the $x$-cycle, in complete agreement with the conclusion reached in the previous section, and the same for $T^{y}$ by choosing $n_{x}={\cal K},n_{y}=1$.
@@ 7.1 Pure TQFT – no charge
In App. B, we will consider a generalization of this.
@@
    $\begin{split}&h(a_{j})={rj^{2}\over 2s}\text{ mod }1\\ &B(a_{j},a_{j^{\prime}})={rjj^{\prime}\over s}\text{ mod }1\,,\end{split}$ (7.6)
@@
The symmetry lines of any spin TQFT can be written as a product of a set of symmetry lines of a bosonic theory and the almost trivial, spin TQFT, $\mathcal{A}^{2,2}\cong\mathcal{A}^{2,0}\cong\{1,c\}$. See App. C.1. In particular, every Abelian spin TQFT is a product of an Abelian bosonic TQFT and $\{1,c\}$.
Let us demonstrate it for the symmetry lines $\mathcal{A}^{s,r}$. If $rs$ is even, $\mathcal{A}^{s,r}$ can be viewed as bosonic, and the general statement above is trivially true. If $rs$ is odd, $\mathcal{A}^{s,r}$ is not bosonic. However, using (7.12), as spin theories, ${\cal A}^{s,r}\cong{\cal A}^{s,r+s}$. Since for odd $rs$, ${\cal A}^{s,r+s}$ also makes sense as a bosonic TQFT, the spin theory $\mathcal{A}^{s,r}$ factorizes as ${\cal A}^{s,r+s}\boxtimes\{1,c\}$, where the first factor is bosonic. Alternatively, using $m=2$ in (7.13), as spin theories, $\mathcal{A}^{s,r}\cong\mathcal{A}^{s,4r}$ Since $\mathcal{A}^{s,4r}$ also makes sense as a bosonic TQFT, the spin theory $\mathcal{A}^{s,r}$ factorizes as $\mathcal{A}^{s,4r}\boxtimes\{1,c\}$, where the first factor is bosonic.
@@
[…]
Since the two labels of ${\cal A}^{{q},{p}}$ are co-prime this means that after gauging the full theory factorizes as
[…]
@@ 7.2 Bosonic TQFT coupled to $\mathrm{U}(1)$
Once we have this group of symmetry lines, we should assign $\mathrm{U}(1)$ charges. The $\mathrm{U}(1)$ charges and the coupling to the background gauge field $A$ depend on a class in $H^{2}(\mathrm{U}(1),\mathbb{Z}_{s})=\mathbb{Z}_{s}$. This class is labeled by an integer $u$ modulo $s$ [^28] As in (7.13), it is sometimes the case that different values of $u$ are related by an automorphism of the TQFT and lead to essentially the same coupling of $A$ to the TQFT. This will be addressed in App. B. Physically, the value of $u$ corresponds to the choice of vison. For simplicity, in the main text, we consider the group of symmetry lines to be generated by the vison, which effectively sets $u=1$. (In App. B we will restore generic values of $u$.)
@@
and all the lines in ${\cal T}^{\prime}$ are $\mathrm{U}(1)$ neutral. (More precisely, we can attach to them the transparent charged boson to set their $\mathrm{U}(1)$ charge to zero.)
@@
In the case of $\ell=\gcd(r,s)\neq 1$, we can follow the discussion around (7.15), (7.16), and (7.17), and try to gauge the $\mathbb{Z}_{\ell}^{(1)}\subset\mathbb{Z}_{s}^{(1)}$ that is generated by $a_{1}^{s/\ell}=v^{q}$. This line generates ${\cal A}^{\ell,{pq\ell}}$. These lines have integer charges and can thus be taken to be neutral. For even $pq$, ${\cal A}^{\ell,{pq\ell}}\cong{\cal A}^{\ell,0}$ and we can gauge it, i.e., it plays the role of ${\cal A}_{0}$ in Sec. 3.2. Here it is important that the lines in ${\cal A}^{\ell,0}$ are $\mathrm{U}(1)$ neutral. Then, as in the problem without $\mathrm{U}(1)$,
@@ 7.3 TQFT coupled to spinc connection
Now we consider coupling our TQFT with symmetry lines ${\cal A}^{s,r}$ to a spinc connection. Recall that, as reviewed in Sec. 3, a spinc theory is a fermionic/spin TQFT satisfying the spin/charge relation. Namely, all local particles with odd charge are identified with the electron, $c$. This relation is often easier to deal with mathematically if we keep the electron as an anyon in our theory, so we make that choice here.
In App. C, we will review the classification of spinc TQFTs in a way that incorporates information about the charges of all of the anyons in the theory. There, we will see that these theories still possess a unique Abelian vison. The braiding of this vison with all other anyons continues to determine their charge modulo one. Moreover, the vison continues to have $(h(v),Q(v))\sim(\sigma_{H}/2,\sigma_{H})$ modulo the identifications discussed in Sec. 3. Thus, if the vison has order $s=q\ell$ it generates ${\cal V}^{s,r}$.
The classification of charge assignments of anyons is more subtle in the spinc case, and we defer a full discussion to App. C. The first new subtlety is that the spin/charge relation introduces additional consistency conditions and therefore the vison cannot be an arbitrary Abelian anyon. The second subtlety is that some superficially distinct choices are actually the same.
@@
This demonstrates the first subtlety above that excludes $Q(a_{1})={1\over 2}$. To demonstrate the second issue, we compare the choices $Q(a_{1})=0\text{ mod }2$ and $Q(a_{1})=1\text{ mod }2$. By changing the generator of $\mathcal{A}^{2,1}$ from $a_{1}$ to $a_{1}c$, the anyon now has vanishing charge, but its spin is shifted by $1/2$ from $1/4$ to $-1/4$ mod 1. In other words, $\mathrm{U}(1)_{2}\times\mathrm{U}(1)_{-1}$ with $Q(a_{1})=1\text{ mod }2$ is equivalent to $\mathrm{U}(1)_{-2}\times\mathrm{U}(1)_{1}$ with $Q(a_{1})=0\text{ mod }2$.
@@
$-\frac{2}{4\pi}b\wedge db+{1\over 4\pi}\hat{b}\wedge\hat{b}-{1\over 2\pi}A\wedge d\hat{b}\,,$ (7.26)

### v3

@@ 5 Charged particles in a magnetic field
The statement that a fractionally quantized Hall conductivity implies the emergence of Abelian anyon excitations with fractional charge is well known. When restated in terms of the higher-form symmetry of the low-energy theory, the conclusion of Sec. 4 might look strange. Where did the one-form symmetry of the low-energy theory come from? An obvious possible answer is that it can be an emergent symmetry, i.e., a symmetry of the IR theory without a precursor in the UV. Although this is certainly a possibility, in this section, we will demonstrate that in the case of particles in a uniform magnetic field, the IR one-form symmetry and its anomaly started their lives in the UV theory as ordinary space translations.
@@ 5.1 The UV theory
$\begin{split}w_{x}&=\exp\left(i\int dx^{\prime}A_{x}^{(0)}(x^{\prime},y)\right)=\exp\left(-2\pi i{y\over L_{y}}\right)\\ w_{y}&=\exp\left(i\int dy^{\prime}A_{y}^{(0)}(x,y)\right)=\exp\left(2\pi i{x\over L_{x}}\right)\,.\end{split}$ (5.4)
@@
The parameters $\sigma_{i}$ correspond to $\delta A_{i}={\sigma_{i}/L_{i}}$. Changing the origin of $x$ or $y$ shifts $\sigma_{y}$ and $\sigma_{x}$ . Hence, these constants signal the fact that the classical $\mathrm{U}(1)$ translation symmetry around every direction is only $\mathbb{Z}_{\cal K}$, rather than $\mathrm{U}(1)$. They are generated by
@@
In the first equation in (5.8), $\sigma_{i}$ are related to a background $\delta A_{i}^{(0)}$ and can be thought of as due to topological defects associated with the global $\mathrm{U}(1)$ symmetry. (Standard flux threading discussions analogous to those in Sec. 2.2 and its counterpart in Sec. 4 shift these values.) The analysis of such defects leads to a mixed anomaly between the discrete ${\mathbb{Z}}_{\cal K}$ translations and $\mathrm{U}(1)$, which in turn leads to the famous filling constraints [85, 86, 87, 88]. Such a change of the background field by $\sigma_{i}$ with vanishing field strength is often referred to as “flux insertion,” and it is done when the total field strength vanishes. Here, the background magnetic field is nonzero, but we can still deform the theory with $\delta A$ with vanishing field strength and use the standard techniques of such defects.
The second equation in (5.8) means that this discrete translation symmetry is realized projectively. This can be interpreted as anomalies in the discrete translation symmetry and the global $\mathrm{U}(1)$ symmetry. [^23] See Appendix A in [89], for a pedagogical discussion of the well-known Landau levels from the perspective of ’t Hooft and Adler-Bell-Jackiw anomalies. This interpretation is significant because it leads to ’t Hooft anomaly matching conditions. In particular, the same anomalous phases should exist in the IR theory. We will use this constraint to classify the possible gapped phases.
@@ 5.2 From the UV to the IR
Let us start with the third question. Most of the symmetry (5.8) is anomaly-free, and therefore it can act trivially in the IR. Let the generators $T^{i}$ flow in the IR to $T_{\rm IR}^{i}$. Then, the elements $(T_{\rm IR}^{i})^{q}$ do not participate in the anomaly and they can act trivially; i.e., act as central elements in the IR theory. We will allow a more general situation depending on another integer of order one $\ell$ and define [^24] This matches with the notation used in Secs. 4 and 7. Comparing with the discussions in Appendix B, here we have $u=1$.
@@
$^{q\over 4\pi}\hat{b}\wedge d\hat{b}-{1\over 2\pi}\delta A\wedge d\hat{b}\,$ (5.14)
@@
where $\oint_{i}$ is a contour integral around the $i$ direction. (For odd $q$ we also need a transparent fermion line on the right-hand side.) This is consistent with the fact that this theory has $p=r=1$.
@@
Finally, given that the IR version of the translation generators $T_{\rm IR}^{i}$ act in the IR as one-form symmetry operators, we can tie this system to the flux threading argument in Secs. 2.2 and 4. We identify $T_{\rm IR}^{i}$ with the one-form symmetry generated by the vison $v$. Then, the first equation in (5.11) is the statement that $v$ carries charge $r/s=p/q$, the second equation in (5.11) is the statement that $B(v,v)=r/s=p/q$, and equation (5.10), which states which subgroup of the translation group is realized trivially in the IR, determines the power of the vison that is trivial. In other words, we have learned that the UV precursor of the vison $v$ is a UV translation symmetry generator.
@@ 6 Connection between translation and flux threading
(b) A cartoon of the action of translation on a TQFT that respects translation and $\mathrm{U}(1)$. (a) Display of the setup in a toroidal geometry with $N_{x}$ unit cells along the $x$ direction and $N_{y}$ along the $y$ direction. The filling constraints ensure that there is a background anyon, $a_{b}$, at the center of each unit cell with a charge $Q(a_{b})=\nu_{l}$. If the system is translated in the $x$ direction by the size of the cell, as in (b), each $a_{b}$ moves to its neighbor. This can be seen to produce $N_{y}$ loop defects of $a_{b}$. In the presence of $2\pi$ flux threading through the hole of the torus, translation thus leads to a phase of $2\pi\nu_{l}N_{y}$. We note that we cannot embed a flat torus in three-dimensional space. Hence, the torus in the figure is not flat. However, since we discuss only discrete translation, we can have interactions preserving them.
@@
We expect a gapped ground state to be described by a TQFT, which might be trivial. Then, from the Lieb-Schulz-Mattis-Oshikawa-Hastings theorem [92, 85, 93, 94], we know that for fractional $\nu_{l}$ gapped ground states must have nontrivial topological order. In particular, they must have an Abelian background anyon, $a_{b}$, with charge $Q(a_{b})=\nu_{l}$. This is because translation acts projectively on every anyon $a$ as $(\hat{T}^{y})^{-1}(\hat{T}^{x})^{-1}\hat{T}^{y}\hat{T}^{x}[a]=e^{2\pi i\varphi(a)}[a]$. [^27] Here $[a]$ refers to the state of a single anyon. In the continuum theory, this is a defect line of $a$ stretching along the time direction. In other words, the action of translation is given by a phase $\varphi$ that respects fusion. (We restrict attention to phases where the action of translation does not permute anyons.) The results of [33], summarized in Appendix C, thus tell us that we can write $(\hat{T}^{y})^{-1}(\hat{T}^{x})^{-1}\hat{T}^{y}\hat{T}^{x}[a]=e^{2\pi iB(a,a_{b})}[a]$ for some Abelian $a_{b}$. The projective action of translation thus means we can think of the background as having $a_{b}$ in each unit cell [95, 96, 97, 86, 98]. As such $a_{b}$ is often called the “background anyon.”
@@
[…]
@@
Then, from our discussion earlier, the action of $\hat{T}^{x}$ corresponds to a line operator of $v$ around the $x$ cycle. Since the unit cell has length $L_{y}/n_{x}$ along the $y$ direction, there are $n_{x}$ defect loops parallel to each other. We see that the action of $\hat{T}^{x}$ corresponds, in the IR theory, to a line operator of $v^{n_{x}}$ along the $x$ cycle. This identification does not depend on the strength of the perturbing potential, and hence holds even in the limit that the potential is turned off. Repeating the same argument for the $y$ direction, we see that $\hat{T}^{y}$ corresponds in the IR to a line defect of $v^{n_{y}}$ along the $y$ cycle.
In Sec. 5, we discussed the magnetic translation symmetry $\mathbb{Z}_{\cal K}\times\mathbb{Z}_{\cal K}$ of particles in a homogeneous magnetic field with its anomaly ${\cal N}\over{\cal K}$. The relation to the $\mathbb{Z}_{n_{x}}\times\mathbb{Z}_{n_{y}}$ translation symmetry discussed here is that the latter is an anomaly free subgroup of the former. For example, we can specialize to $n_{x}=1,n_{y}={\cal K}$, which leads to the conclusion that $\hat{T}^{x}=T^{x}$ creates a line defect of $v$ along the $x$ cycle, in complete agreement with the conclusion reached in the previous section, and the same for $T^{y}$ by choosing $n_{x}={\cal K},n_{y}=1$.
@@ 7.1 Pure TQFT – no charge
In Appendix B, we will consider a generalization of this.
@@
[…]
@@
The symmetry lines of any spin TQFT can be written as a product of a set of symmetry lines of a bosonic theory and the almost trivial, spin TQFT, $\mathcal{A}^{2,2}\cong\mathcal{A}^{2,0}\cong\{1,c\}$. See Appendix C.1. In particular, every Abelian spin TQFT is a product of an Abelian bosonic TQFT and $\{1,c\}$.
Let us demonstrate it for the symmetry lines $\mathcal{A}^{s,r}$. If $rs$ is even, $\mathcal{A}^{s,r}$ can be viewed as bosonic, and the general statement above is trivially true. If $rs$ is odd, $\mathcal{A}^{s,r}$ is not bosonic. However, using (7.12), as spin theories, ${\cal A}^{s,r}\cong{\cal A}^{s,r+s}$. Since for odd $rs$, ${\cal A}^{s,r+s}$ also makes sense as a bosonic TQFT, the spin theory $\mathcal{A}^{s,r}$ factorizes as ${\cal A}^{s,r+s}\boxtimes\{1,c\}$, where the first factor is bosonic. Alternatively, using $m=2$ in (7.13), as spin theories, $\mathcal{A}^{s,r}\cong\mathcal{A}^{s,4r}$ . Since $\mathcal{A}^{s,4r}$ also makes sense as a bosonic TQFT, the spin theory $\mathcal{A}^{s,r}$ factorizes as $\mathcal{A}^{s,4r}\boxtimes\{1,c\}$, where the first factor is bosonic.
@@
${{\cal A}^{s,r}\over{\cal A}^{\ell,0}}\cong{\cal A}^{{q},{p}}\quad,\quad{\rm bosonic}\quad pq\in 2\mathbb{Z}\qquad\text{or fermionic/spin}\,.$ (7.15)
Since the two labels of ${\cal A}^{{q},{p}}$ are coprime this means that after gauging the full theory factorizes as
${{\cal T}\over{\cal A}^{\ell,0}}={\cal A}^{{q},{p}}\boxtimes{\cal T}^{\prime}\quad,\quad{\rm bosonic}\quad pq\in 2\mathbb{Z}\qquad\text{or fermionic/spin}\,.$ (7.16)
@@ 7.2 Bosonic TQFT coupled to $\mathrm{U}(1)$
Once we have this group of symmetry lines, we should assign $\mathrm{U}(1)$ charges. The $\mathrm{U}(1)$ charges and the coupling to the background gauge field $A$ depend on a class in $H^{2}(\mathrm{U}(1),\mathbb{Z}_{s})=\mathbb{Z}_{s}$. This class is labeled by an integer $u$ modulo ss. [^28] As in (7.13), it is sometimes the case that different values of $u$ are related by an automorphism of the TQFT and lead to essentially the same coupling of $A$ to the TQFT. This will be addressed in Appendix B. Physically, the value of $u$ corresponds to the choice of vison. For simplicity, in the main text, we consider the group of symmetry lines to be generated by the vison, which effectively sets $u=1$. (In Appendix B we will restore generic values of $u$.)
@@
and all the lines in $\cal T$$^{\prime}$ are $\mathrm{U}(1)$ neutral. [More precisely, we can attach to them the transparent charged boson to set their $\mathrm{U}(1)$ charge to zero.]
@@
In the case of $\ell=\gcd(r,s)\neq 1$, we can follow the discussion around (7.15)-(7.17), and try to gauge the $\mathbb{Z}_{\ell}^{(1)}\subset\mathbb{Z}_{s}^{(1)}$ that is generated by $a_{1}^{s/\ell}=v^{q}$. This line generates ${\cal A}^{\ell,{pq\ell}}$. These lines have integer charges and can thus be taken to be neutral. For even $pq$, ${\cal A}^{\ell,{pq\ell}}\cong{\cal A}^{\ell,0}$ and we can gauge it, i.e., it plays the role of ${\cal A}_{0}$ in Sec. 3.2. Here it is important that the lines in ${\cal A}^{\ell,0}$ are $\mathrm{U}(1)$ neutral. Then, as in the problem without $\mathrm{U}(1)$,
@@ 7.3 TQFT coupled to spinc connection
Now we consider coupling our TQFT with symmetry lines ${\cal A}^{s,r}$ to a spinc connection. Recall that, as reviewed in Sec. 3, a spinc theory is a fermionic/spin TQFT satisfying the spin/charge relation. Namely, all local particles with odd charge are identified with the electron $c$. This relation is often easier to deal with mathematically if we keep the electron as an anyon in our theory, so we make that choice here.
In Appendix C, we will review the classification of spinc TQFTs in a way that incorporates information about the charges of all of the anyons in the theory. There, we will see that these theories still possess a unique Abelian vison. The braiding of this vison with all other anyons continues to determine their charge modulo one. Moreover, the vison continues to have $(h(v),Q(v))\sim(\sigma_{H}/2,\sigma_{H})$ modulo the identifications discussed in Sec. 3. Thus, if the vison has order $s=q\ell$ it generates ${\cal V}^{s,r}$.
The classification of charge assignments of anyons is more subtle in the spinc case, and we defer a full discussion to Appendix C. The first new subtlety is that the spin/charge relation introduces additional consistency conditions and therefore the vison cannot be an arbitrary Abelian anyon. The second subtlety is that some superficially distinct choices are actually the same.
@@
This demonstrates the first subtlety above that excludes $Q(a_{1})={1\over 2}$. To demonstrate the second issue, we compare the choices $Q(a_{1})=0\text{ mod }2$ and $Q(a_{1})=1\text{ mod }2$. By changing the generator of $\mathcal{A}^{2,1}$ from $a_{1}$ to $a_{1}c$, the anyon now has vanishing charge, but its spin is shifted by $1/2$ from $1/4$ to 1/4-1/4 mod 1. In other words, $\mathrm{U}(1)_{2}\times\mathrm{U}(1)_{-1}$ with $Q(a_{1})=1\text{ mod }2$ is equivalent to $\mathrm{U}(1)_{-2}\times\mathrm{U}(1)_{1}$ with $Q(a_{1})=0\text{ mod }2$.
@@
$\frac{2}{4\pi}b\wedge db+{1\over 4\pi}\hat{b}\wedge\hat{b}-{1\over 2\pi}A\wedge d\hat{b}\,$, (7.26)